a candy manufacturer has two frades of candy selling at Php100/kilo and Php120/kilo respectively. how many kilos of each must be used to make a mixture of 50 kilos, to sell at Php105/ kilo?
please help me .

Answer by mananth(16946)

Php ---------------- quantity
Candy type I 100 ---------------- x Kg
candy type II 120 ------ 50 - x Kg
Mixture 105 ---------------- 50
600
100 x + 120 ( 50 - x ) = 105 * 50
10000 x + 12000 ( 50 - x ) = 525000
10000 x + 600000 - 12000 x = 525000
10000 x - 12000 x = 525000 - -600000
-2000 x = -75000
/ -2000
x = 37.5 Kg Php100.00 Candy type I
12.5 Kg Php 120.00 candy type II

m.ananth@hotmail.ca